# Efficient Monte-Carlo sampling of metastable systems using non-local collective variable updates

### v2

@@ B.2.4 Symplecticity and reversibility of the deterministic flows
Consider the the numerical flow
@@
To prove this statement, we show that the 2-form $\mathrm{d}q\wedge\mathrm{d}p$ is preserved, where $\wedge$ denotes the wedge-product (see Ref. McDuffSalamonSymplecticTop for a textbook introduction to symplectic structures). For the simplest version with a constant constraint, this is a standard result found for example in Refs. hairer-lubich-wanner-06 and leimkuhler_symplectic_1994. We follow here the same arguments as in Ref. leimkuhler_symplectic_1994. The differentials obey:
@@
As in the proof by Ref. leimkuhler_symplectic_1994, we make use of the following two lemmata.
@@
In all of the following, we assume that the time-step $\Delta t$ is sufficiently small so that, along the trajectory, the forward and backward flow associated with the deterministic parts (II.18) and (B.10) satisfy a reversibility property, see e.g. Sect. VII.1.4 of Ref. hairer-lubich-wanner-06 and Sect. 5.4.3 of Ref. lelievre_langevin_2012, as well as Lemma 2.9 in Ref. lelievre-rousset-stoltz-19 where a precise setting is exhibited for which this is satisfied. More precisely, we assume that the following pointwise equality holds for the forward and backward flows $\Phi^{k}$ and $\Phi^{{\rm b},k^{\prime}}$ introduced in (B.14) and (B.15). For all the configurations which are visited and all $k\in\{1,\ldots,K_{T}\}$:
@@ B.2.5 Jarzynski–Crooks equality
We are now in position to state the Jarzynski–Crooks formula. We state it here in a slightly more general form than in Ref. lelievre_langevin_2012, but otherwise follow exactly the presentation there.
@@
The proof is the same as the proof of Ref. lelievre_langevin_2012, Theorem 5.5. The only difference lies in the slightly different momentum constraint $(C_{p})$ for the forward and backward dynamics. Since we show in ˜B.6 that the forward and backward deterministic flows are symplectic, the proof goes through in the same way. ∎
@@ B.3 Proof of reversibility
We are now in a position to prove Theorem˜II.4, namely that the algorithm presented in Section˜II.2 is reversible with respect to the target probability measure $\nu$. Assume that $Q_{n}$ is distributed according to $\nu$. The aim is to prove that $(Q_{n},Q_{n+1})$ has the same law as $(Q_{n+1},Q_{n})$. In order to study the law of $(Q_{n},Q_{n+1})$, let us consider a bounded measurable function $\varphi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. Recall that $Z_{n}=\xi(Q_{n})$ and $\widetilde{Z}_{n+1}=\xi(\widetilde{Q}_{n+1})$. One has:
@@
    $\nu(\mathrm{d}q\,|\,\xi(q)=z)=\frac{Z_{q}^{-1}e^{-\beta V(q)}\delta_{\xi(q)-z}(\mathrm{d}q)}{Z(z)},$
    where $Z(z)=Z_{q}^{-1}\int_{\Sigma(z)}e^{-\beta V(q)}\delta_{\xi(q)-z}(\mathrm{d}q)=\exp(-\beta F(z))$ is the normalizing constant.
@@
We can use Eq.˜B.13 to rewrite the latter expectation value over a backward path as one over a forward path as
@@
which concludes the proof of Theorem˜II.4.
@@ Appendix D Simplification for a linear CV
To explicitly connect the presented algorithm to the previous one from Ref. SCHONLE2025113806, consider the simplest case where the reaction coordinate is just given by the first $\ell$ degrees of freedom, so that we can write $q=(q_{\mathrm{CV}},q_{\perp})$ with $q_{\mathrm{CV}}=\xi(q)=(q_{1},\dots,q_{\ell})$ and $q_{\perp}=(q_{\ell+1},\dots,q_{d})$. Similarly, one can then write the momenta as $p=(p_{\mathrm{CV}},p_{\perp})$. Assume once again that $M=M\operatorname{Id}_{d}$ (with a slight abuse of notation). Then the Gram tensor is $G_{M}(q)=M^{-1}\operatorname{Id}_{\ell}$, and the projection operator for the momenta is just
@@ D.1 Overdamped limit
In terms of the parameters of the normalized algorithm introduce in Section˜II.4, this parameter choice corresponds to setting $\alpha_{1}=1$. Using (II.22), the associated work is calculated as
@@
This algorithm is very closely related to the asymmetric algorithm from Ref. SCHONLE2025113806, with the only difference that forward and backward path have swapped their definitions.
@@ D.2 Deterministic limit
The last equality, where the norm of the momentum is only considered on the orthogonal degrees of freedom, comes from the assumption that the initial and final velocities in CV space are equal. This is the relaxed version of (II.6) when there is no Fixman term to be considered, see Remark˜II.3.
@@ Appendix E Details of the wavelet basis for the $\phi^{4}$ model
As mentioned in Section˜III.2, we use a basis of Haar-wavelets to represent the state of the $\phi^{4}$ model in our simulation. For an introduction to wavelet transformations we refer the reader to Refs. chui1992; mallat_wavelet_2008, for example. Here, we only give a very practical definition. For a one dimensional input signal $\{\phi_{i}\}_{i=1}^{N}$ of size $N=2^{n}$, set $\varphi^{(0)}_{i}=\phi_{i}$ for $i\in\{1,\ldots N\}$ and repeatedly apply the following transformation $n$ times for $j\in\{1,\ldots n-1\}$:
@@ F.2 Lagrange multiplier for position constraint
We choose the smallest value for $\lambda^{k+1/2}$, which corresponds to choosing $\lambda^{k+1/2}_{+}$ if $e_{12}^{k}\cdot(\tilde{q}^{k+1}_{1}-\tilde{q}^{k+1}_{2})>0$ and $\lambda^{k+1/2}_{-}$ in the opposite case. This choice is consistent with ensuring that the value of the Lagrange multiplier is of order $\Delta t$ for small timesteps.
@@ Appendix G Details of the polymer model
We simulate the polymer system introduced in Section˜III.4 at a temperature $1/\beta=1.3$ using the jaxMD packagejaxmd2020. The system is contained in a three-dimensional box of length $L=7.47$ with periodic boundaries, with $N_{\mathrm{solvent}}=200$ solvent particles, which yields a solvent density of $N_{\mathrm{solvent}}/L^{3}=0.48$ particles per unit volume. The parameters of the polymer potentials are $d_{1}=1$, $d_{2}=1.63$ (to fix the bonding angle at 109°), $k_{1}=k_{2}=500$, $d_{e}=4.62$, $a_{e}=0.8$ and $k_{e}=9.1$, corresponding to an energy barrier of $7\beta^{-1}$. The Lennard-Jones potential is essentially given by
@@
To implement Step 1 of the algorithm in Section˜II.2, we use an independent proposal sampler in CV space for both choices of CV considered here. For the first CV, i.e. the one-dimensional end-to-end distance, the proposal sampler is given by a Gaussian mixture model with two modes of equal weight and width $\sigma=0.17$ at positions $z=3.53$ and $z=5.65$. For the second, 27-dimensional CV, i.e. the Cartesian coordinates of the polymer beads, we use a normalizing flow to propose distances between the polymer beads. These distances are the eight nearest neighbor and seven second-nearest neighbor distances as well as the six distances between the first bead and beads number 4 to 9, adding up to 21 distances in total. To complete such a proposed set of distances to a full proposal of polymer coordinates, one needs to additionally specify the global translation and rotation, as well as six sign values $\eta_{i}\in\{\pm 1\},\,i\in\{1,\ldots 6\}$ (explained in the following). The global translation is fixed by always positioning the central bead of the polymer at the center of the box (which we do throughout the simulation) and the global rotation is sampled uniformly. The necessity of the six sign values $\eta_{i}$ can be understood from the triangulation process used to determine the Cartesian polymer bead positions. Assuming one has already placed three beads with the desired distances, there will be two possible solutions (if there is a solution) to place the fourth bead at the right distance to the previous three (two intersection points of three spheres), which corresponds to choosing the sign of the dihedral angle. For a given set of 21 distances as in our case, there are six such sign values to choose, and we sample them uniformly at random. We compute the Jacobian of this reconstruction using automatic differentiation to arrive at a probability measure in the CV space of polymer positions $z\in\mathbb{R}^{27}$. The flow will sometimes produce sets of distances for which no solution to the triangulation problem exists and these samples are then directly rejected. Since this restriction only changes the normalization of the flow proposal measure, it does not enter into the overall formula of the accept/reject probability in Step 3 of the algorithm in Section˜II.2. The flow is implemented using the FlowJax packageward2023flowjax, with one affine whitening layer and 20 blocks of RationalQuadraticSpline transformationsdurkan2019neuralsplineflows with 32 knots on the interval [4,4][-4,4] and internal neural networks of depth 2 and width 30 (see the reference for further explanations). It was trained on a large dataset of one million samples obtained from a long Langevin simulation over 200 epochs with a batch size of 40,000 and a learning rate of $10^{-3}$ using the Adam optimizer with otherwise default parameters.
@@ H.1 Gaussian tunnel
To illustrate the optimal choice of parameters when running the algorithm on the example of the Gaussian tunnel from Section˜III.1, we show the inverse mode jump cost (introduced in the main text) for different parameter values in Fig.˜13. The optimal value of $\alpha_{2}$ is roughly the same for all choices of $\alpha_{1}$, whereas the optimal number of intermediate steps $K_{T}$ increases with $\alpha_{1}$. Clearly, the optimal performance is observed for the deterministic dynamics, namely $\alpha_{1}=0$.
@@
For the deterministic case, there is a simple continuous-time limit of $\Delta t\longrightarrow 0$, in our parameterization corresponding to $\alpha_{2}\longrightarrow 0$, where the only relevant quantity determining the acceptance rate is the normalized transition time $T/\sqrt{\beta M}$. This is indeed what we observe in Fig.˜14, where for small enough $\alpha_{2}$, the acceptance rate only depends on this physical time scale. For numerical performance, however, the relevant quantity is of course the acceptance rate normalized by the numerical cost.
Figure 14: Acceptance rate for a fixed jump from $z=0$ to $\widetilde{z}=b$ for the Gaussian tunnel with the deterministic dynamics $\alpha_{1}=0$. The results shown here are the same as in the top left panel of Fig.˜1 with a different choice of axes. The right panel shows all simulations with $\alpha_{2}<0.4$ combined in one plot.
@@ H.2 Gaussian tunnel with a non-linear collective variable
We consider here a version of the Gaussian tunnel introduced in Section˜III.1 in dimension $d=10$ with a non-linear collective variable. With the same underlying probability distribution, we consider the CV $\xi(z)=\tanh\left(\frac{z}{b}\right)\cdot\frac{b}{\tanh(1)}$. Using the push-forward of the probability density $\nu_{\mathrm{CV}}(z)$ under this transformation as the proposal distribution (and thus exactly the marginal of the target distribution), we show in Fig.˜15 how our algorithm with the schedule (II.11) leads to unbiased results, whereas a linear schedule that violates Eq.˜II.6 or the omission of the Fixman term both lead to a bias. To test an inappropriate schedule, we simply use one with constant velocity, i.e. (II.9) and (II.10) with $f(\tau)=\tau$. The Fixman term is given by $V_{\mathrm{fix}}(z,x^{\perp})=\frac{1}{\beta}\log(\xi^{\prime}(z))$ and just depends on $z$. We observe that upon omission of the Fixman term, the resulting distribution is indeed biased: its marginal along $\xi$ is proportional to $\nu_{\mathrm{CV}}(z)\cdot e^{\beta V_{\mathrm{fix}}(z)}$.
@@ H.3 $\phi^{4}$ Model
To complement the results shown in Fig.˜4 we ran the algorithm on the $\phi^{4}$ model from Section˜III.2 for additional values of $\alpha_{1}$ and show the associated mode-jump cost in Fig.˜16.

### v3

@@ B.2.4 Symplecticity and reversibility of the deterministic flows
Consider the numerical flow
@@
To prove this statement, we show that the 2-form $\mathrm{d}q\wedge\mathrm{d}p$ is preserved, where $\wedge$ denotes the wedge-product (see Ref. McDuff and Salamon, 2017 for a textbook introduction to symplectic structures). For the simplest version with a constant constraint, this is a standard result found for example in Refs. Hairer et al., 2006 and Leimkuhler and Skeel, 1994. We follow here the same arguments as in Ref. Leimkuhler and Skeel, 1994. The differentials obey:
@@
As in the proof by Ref. Leimkuhler and Skeel, 1994, we make use of the following two lemmata.
@@
We now make more precise the reversibility property we require between the forward and backward flow, already mentioned in Remark˜B.3. In all of the following, we assume that the time-step $\Delta t$ is sufficiently small so that, along the trajectory, the forward and backward flows $\Phi^{k}$ and $\Phi^{{\rm b},k^{\prime}}$ introduced in (B.14) and (B.15) and associated with the deterministic parts (II.18) and (B.10) satisfy the following property: For all the configurations which are visited and all $k\in\{1,\ldots,K_{T}\}$ the following pointwise equality holds:
@@
See e.g. Sect. VII.1.4 of Ref. Hairer et al., 2006 and Sect. 5.4.3 of Ref. Lelièvre et al., 2012, as well as Lemma 2.9 in Ref. Lelièvre et al., 2019 where a precise setting for which this is satisfied is exhibited. Potential difficulties around this could also be dealt with using so-called reversibility checks, but we leave this for future work.
@@ B.2.5 Jarzynski–Crooks equality
We are now in position to state the Jarzynski–Crooks formula. We state it here in a slightly more general form than in Ref. Lelièvre et al., 2012, but otherwise follow exactly the presentation there.
@@
The proof is the same as the proof of Ref. Lelièvre et al., 2012, Theorem 5.5 and relies in particular on the reversibility property of the forward and backward dynamics (B.21). The only difference lies in the slightly different momentum constraint $(C_{p})$ for the forward and backward dynamics. Since we show in ˜B.6 that the forward and backward deterministic flows are symplectic, the proof goes through in the same way. ∎
@@ B.3 Proof of reversibility
We are now in a position to prove Theorem˜II.3, namely that the algorithm presented in Section˜II.2 is reversible with respect to the target probability measure $\nu$. Assume that $Q_{n}$ is distributed according to $\nu$. The aim is to prove that $(Q_{n},Q_{n+1})$ has the same law as $(Q_{n+1},Q_{n})$. In order to study the law of $(Q_{n},Q_{n+1})$, let us consider a bounded measurable function $\varphi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. Recall that $Z_{n}=\xi(Q_{n})$ and $\widetilde{Z}_{n+1}=\xi(\widetilde{Q}_{n+1})$. One has:
@@
    $\nu(\mathrm{d}q\,|\,\xi(q)=z)=\frac{\mathcal{Z}_{q}^{-1}e^{-\beta V(q)}\delta_{\xi(q)-z}(\mathrm{d}q)}{\exp(-\beta F(z))},$
    where $\exp(-\beta F(z))$ is the normalizing constant, recall (B.1).
@@
We can use ˜B.5 to rewrite the latter expectation value over a backward path as one over a forward path as
@@
which concludes the proof of Theorem˜II.3.
@@ Appendix D Simplification for a linear CV
To explicitly connect the presented algorithm to the previous one from Ref. Schönle et al., 2025, consider the simplest case where the reaction coordinate is just given by the first $\ell$ degrees of freedom, so that we can write $q=(q_{\mathrm{CV}},q_{\perp})$ with $q_{\mathrm{CV}}=\xi(q)=(q_{1},\dots,q_{\ell})$ and $q_{\perp}=(q_{\ell+1},\dots,q_{d})$. Similarly, one can then write the momenta as $p=(p_{\mathrm{CV}},p_{\perp})$. Assume once again that $M=M\operatorname{Id}_{d}$ (with a slight abuse of notation). Then the Gram tensor is $G_{M}(q)=M^{-1}\operatorname{Id}_{\ell}$, and the projection operator for the momenta is just
@@
This corresponds to the underdamped Langevin dynamics on $(q_{\perp},p_{\perp})$ with a time-varying potential $V_{k}(q^{k}_{\perp})$, which becomes obvious by rewriting $V(q^{k})=V(q^{k}_{\mathrm{CV}},q^{k}_{\perp})\equiv V_{k}(q^{k}_{\perp})$, since the CV-space coordinates $q_{\mathrm{CV}}$ are fully determined by the schedule. One might therefore also see the CV coordinates $q_{\mathrm{CV}}^{k}$ here as a time-varying alchemical parameter.
@@ D.1 Overdamped limit
In terms of the parameters of the normalized algorithm introduced in Section˜II.4, this parameter choice corresponds to setting $\alpha_{1}=1$. Using (II.22), the associated work is calculated as
@@
This algorithm is very closely related to the asymmetric algorithm from Ref. Schönle et al., 2025, with the only difference that forward and backward path have swapped their definitions.
@@ D.2 Deterministic limit
The last equality, where the norm of the momentum is only considered on the orthogonal degrees of freedom, comes from the assumption that the initial and final velocities in CV space are equal. This is the relaxed version of (II.6) when there is no Fixman term to be considered, see Remark˜II.2.
@@ Appendix E Details of the wavelet basis for the $\phi^{4}$ model
As mentioned in Section˜III.2, we use a basis of Haar-wavelets to represent the state of the $\phi^{4}$ model in our simulation. For an introduction to wavelet transformations we refer the reader to Refs. Chui, 1992; Mallat, 2008, for example. Here, we only give a very practical definition. For a one-dimensional input signal $\{\phi_{i}\}_{i=1}^{N}$ of size $N=2^{n}$, set $\varphi^{(0)}_{i}=\phi_{i}$ for $i\in\{1,\ldots N\}$ and repeatedly apply the following transformation $n$ times for $j\in\{1,\ldots n-1\}$:
@@ F.2 Lagrange multiplier for position constraint
We choose the smallest value for $\lambda^{k+1/2}$, which corresponds to choosing $\lambda^{k+1/2}_{+}$ if $e_{12}^{k}\cdot(\tilde{q}^{k+1}_{1}-\tilde{q}^{k+1}_{2})>0$ and $\lambda^{k+1/2}_{-}$ in the opposite case. This choice is consistent with ensuring that the value of the Lagrange multiplier is of order $\Delta t$ for small timesteps and ensures that (B.21) is satisfied.
@@ Appendix G Details of the polymer model
We simulate the polymer system introduced in Section˜III.4 at a temperature $1/\beta=1.3$ using the jaxMD packageSchoenholz and Cubuk (2020). The system is contained in a three-dimensional box of length $L=7.47$ with periodic boundaries, with $N_{\mathrm{solvent}}=200$ solvent particles, which yields a solvent density of $N_{\mathrm{solvent}}/L^{3}=0.48$ particles per unit volume. The parameters of the polymer potentials are $d_{1}=1$, $d_{2}=1.63$ (to fix the bonding angle at 109°), $k_{1}=k_{2}=500$, $d_{e}=4.62$, $a_{e}=0.8$ and $k_{e}=9.1$, corresponding to an energy barrier of $7\beta^{-1}$. The Lennard-Jones potential is essentially given by
@@
To implement Step 1 of the algorithm in Section˜II.2, we use an independent proposal sampler in CV space for both choices of CV considered here. For the first CV, i.e. the one-dimensional end-to-end distance, the proposal sampler is given by a Gaussian mixture model with two modes of equal weight and width $\sigma=0.17$ at positions $z=3.53$ and $z=5.65$. For the second, 27-dimensional CV, i.e. the Cartesian coordinates of the polymer beads, we use a normalizing flow to propose distances between the polymer beads. These distances are the eight nearest neighbor and seven second-nearest neighbor distances as well as the six distances between the first bead and beads number 4 to 9, adding up to 21 distances in total. To complete such a proposed set of distances to a full proposal of polymer coordinates, one needs to additionally specify the global translation and rotation, as well as six sign values $\eta_{i}\in\{\pm 1\},\,i\in\{1,\ldots 6\}$ (explained in the following). The global translation is fixed by always positioning the central bead of the polymer at the center of the box (which we do throughout the simulation) and the global rotation is sampled uniformly. The necessity of the six sign values $\eta_{i}$ can be understood from the triangulation process used to determine the Cartesian polymer bead positions. Assuming one has already placed three beads with the desired distances, there will be two possible solutions (if there is a solution) to place the fourth bead at the right distance to the previous three (two intersection points of three spheres), which corresponds to choosing the sign of the dihedral angle. For a given set of 21 distances as in our case, there are six such sign values to choose, and we sample them uniformly at random. We compute the Jacobian of this reconstruction using automatic differentiation to arrive at a probability measure in the CV space of polymer positions $z\in\mathbb{R}^{27}$. The flow will sometimes produce sets of distances for which no solution to the triangulation problem exists and these samples are then directly rejected. Since this restriction only changes the normalization of the flow proposal measure, it does not enter into the overall formula of the accept/reject probability in Step 3 of the algorithm in Section˜II.2. The flow is implemented using the FlowJax packageWard (2025), with one affine whitening layer and 20 blocks of RationalQuadraticSpline transformationsDurkan et al. (2019) with 32 knots on the interval [4,4][-4,4] and internal neural networks of depth 2 and width 30 (see the reference for further explanations). It was trained on a large dataset of one million samples obtained from a long Langevin simulation over 200 epochs with a batch size of 40,000 and a learning rate of $10^{-3}$ using the Adam optimizer with otherwise default parameters.
@@ H.1 Gaussian tunnel
To illustrate the optimal choice of parameters when running the algorithm on the example of the Gaussian tunnel from Section˜III.1, we show the inverse mode jump cost (introduced in the main text) for different parameter values in Fig.˜14. The optimal value of $\alpha_{2}$ is roughly the same for all choices of $\alpha_{1}$, whereas the optimal number of intermediate steps $K_{T}$ increases with $\alpha_{1}$. Clearly, the optimal performance is observed for the deterministic dynamics, namely $\alpha_{1}=0$.
@@
For the deterministic case, there is a simple continuous-time limit of $\Delta t\longrightarrow 0$, in our parameterization corresponding to $\alpha_{2}\longrightarrow 0$, where the only relevant quantity determining the acceptance rate is the normalized transition time $T/\sqrt{\beta M}$. This is indeed what we observe in Fig.˜15, where for small enough $\alpha_{2}$, the acceptance rate only depends on this physical time scale. For numerical performance, however, the relevant quantity is of course the acceptance rate normalized by the numerical cost.
Figure 15: Acceptance rate for a fixed jump from $z=0$ to $\widetilde{z}=b$ for the Gaussian tunnel with the deterministic dynamics $\alpha_{1}=0$. The results shown here are the same as in the top left panel of Fig.˜2 with a different choice of axes. The right panel shows all simulations with $\alpha_{2}<0.4$ combined in one plot.
@@ H.2 Gaussian tunnel with a non-linear collective variable
We consider here a version of the Gaussian tunnel introduced in Section˜III.1 in dimension $d=10$ with a non-linear collective variable. With the same underlying probability distribution, we consider the CV $\xi(z)=\tanh\left(\frac{z}{b}\right)\cdot\frac{b}{\tanh(1)}$. Using the push-forward of the probability density $\nu_{\mathrm{CV}}(z)$ under this transformation as the proposal distribution (and thus exactly the marginal of the target distribution), we show in Fig.˜16 how our algorithm with the schedule (II.11) leads to unbiased results, whereas a linear schedule that violates Eq.˜II.6 or the omission of the Fixman term both lead to a bias. To test an inappropriate schedule, we simply use one with constant velocity, i.e. (II.9) and (II.10) with $f(\tau)=\tau$. The Fixman term is given by $V_{\mathrm{fix}}(z,x^{\perp})=\frac{1}{\beta}\log(\xi^{\prime}(z))$ and just depends on $z$. We observe that upon omission of the Fixman term, the resulting distribution is indeed biased: its marginal along $\xi$ is proportional to $\nu_{\mathrm{CV}}(z)\cdot e^{\beta V_{\mathrm{fix}}(z)}$.
@@ H.3 $\phi^{4}$ Model
To complement the results shown in Fig.˜5 we ran the algorithm on the $\phi^{4}$ model from Section˜III.2 for additional values of $\alpha_{1}$ and show the associated mode-jump cost in Fig.˜17.